# Conceptualization, Methodology, Investigation, Data curation, Formal analysis, Writing - Original draft preparation

### v4

@@ 7 Conclusions and Future Work
In our study, we comprehensively analyzed the practicality of current popular LLMs in the context of security code review. We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews. For the best-performing LLM, we conducted a linguistic analysis of the quality problems in the responses generated by this LLM. We performed a regression analysis to explore factors influencing its performance.
Our main findings are: (1) In security code review tasks, LLMs demonstrate better capabilities compared to state-of-the-art static analysis tools. However, their application remains limited. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) During code review, GPT-4 is more effective in identifying security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project. (4) The most prevalent quality problems in the responses generated by GPT-4 are verbosity and non-compliance with task requirements.
We plan to extend this work in the next step by incorporating code reviews from proprietary software projects into our dataset and attempting enhancement strategies such as few-shot learning and fine-tuning. In addition to applying LLMs to help reviewers detect security defects during code review, we also plan to analyze the ability of LLMs to fix security defects in code so that developers can comprehensively evaluate the potential of LLMs in the task of secure code review.
@@
We have shared the link to our dataset in the reference (Yu et al., 2025).
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.

### v5

@@ 7 Conclusions and Future Work
In our study, we comprehensively analyzed the practicality of current popular LLMs in the context of security code review. We investigated the capabilities of 7 LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews. For the two top-performing LLMs, each paired with their respective optimal prompts, we conducted a linguistic analysis of the quality problems in their generated responses. Furthermore, we performed separate regression analyses to explore the factors influencing the performance of each LLM-prompt combination.
Our main findings are: (1) In security code review tasks, LLMs demonstrate stronger capabilities than state-of-the-art static analysis tools. Among these LLMs, the reasoning-optimized model outperforms established general-purpose models. However, their practical application remains limited. (2) DeepSeek-R1 is the top-performing LLM in security code review, followed by GPT-4. DeepSeek-R1 performs best when the prompt includes the commit message and CoT guidance, whereas GPT-4 performs best when provided with a CWE list in the prompt. (3) DeepSeek-R1 and GPT-4 are more effective at detecting security defects in files with fewer tokens and with security-related code annotations. Notably, DeepSeek-R1 was found to be less focused on Race Condition than other types of security defect. (4) The most prevalent quality problem for both DeepSeek-R1 and GPT-4 is verbosity. In comparison, GPT-4 tends to produce vague expressions and demonstrates poorer instruction-following ability, while DeepSeek-R1 is more prone to errors in recalling code details.
In future work, we plan to explore several directions to further enhance the capabilities of LLMs in security code review: (1) Improving the sensitivity to security defects of LLMs, by investigating the optimal granularity of the CWE list provided in the prompt and incorporating external knowledge or project-specific context to ensure comprehensive consideration of frequent security defect types; (2) Enhancing contextual reasoning by developing approaches to provide more precise code context to inform the analysis procedure; (3) Increasing the stability of the LLMs’ outputs via techniques such as self-refinement and voting mechanisms, to generate more consistent and reliable results.
@@
We have shared the link to our dataset in the reference (replpack).
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.